OpenStudy (anonymous):

What percent of 60 is 12?

OpenStudy (zzr0ck3r):

x*60 = 12 solve for x then turn it into a decimal

Parth (parthkohli):

Another way to think about it is that 12 is a fifth of 60. What is a fifth of hundred?

OpenStudy (zzr0ck3r):

then turn it into a percent*

OpenStudy (zzr0ck3r):

yes this is also a good idea, but if the numbers are strange use the first method.

OpenStudy (zzr0ck3r):

like what percent of pi is 3.078

Parth (parthkohli):

Yeah.

OpenStudy (mathstudent55):

In maths, "of" ususally means times. Is means equal. Percent means /100 An unknown is x. Put it all together and you get: What percent of 60 is 12? x% of 60 = 12 x/100 * 60 = 12 x/100 = 12/60 x = 12*100/60 x = 20 Answer 20%